### v1

@@ Abstract
The accelerated progress of artificial intelligence (AI) has popularized deep learning models across domains, yet their inherent opacity poses challenges, notably in critical fields like healthcare, medicine and the geosciences. Explainable AI (XAI) has emerged to shed light on these ”black box” models, helping decipher their decision making process. Nevertheless, different XAI methods yield highly different explanations. This inter-method variability increases uncertainty and lowers trust in deep networks’ predictions. In this study, for the first time, we propose a novel framework designed to enhance the explainability of deep networks, by maximizing both the accuracy and the comprehensibility of the explanations. Our framework integrates various explanations from established XAI methods and employs a non-linear ”explanation optimizer” to construct a unique and optimal explanation. Through experiments on multi-class and binary classification tasks in 2D object and 3D neuroscience imaging, we validate the efficacy of our approach. Our explanation optimizer achieved superior faithfulness scores, averaging 155% and 63% higher than the best performing XAI method in the 3D and 2D applications, respectively. Additionally, our approach yielded lower complexity, increasing comprehensibility. Our results suggest that optimal explanations based on specific criteria are derivable and address the issue of inter-method variability in the current XAI literature.
@@
XAI, neuroscience, brain, 3D, 2D, computer vision, classificationMain
@@
In the last decade, deep learning models have led to transformative changes across diverse domains, including the geosciences [1, 2], automotive engineering, healthcare [3, 4, 5], and biology [6, 7]. However, the inherent opacity of these sophisticated models, often likened to ”black boxes,” presents formidable challenges in comprehending their decision-making processes. In response to this challenge, the field of eXplainable Artificial Intelligence (XAI) has emerged as a pivotal area of research dedicated to explaining these black-box models and elucidating the rationale behind their predictions. Over the years, a multitude of XAI methods have been developed, all aimed at shedding light on how artificial intelligence (AI) arrives at specific predictions.
Explainable artificial intelligence can be broadly classified into two methodological approaches: inherently interpretable and post-hoc. Interpretable methods prioritize models endowed with properties such as simulatability, decomposability, and transparency, making them inherently comprehensible. Typically, these include linear techniques like Bayesian classifiers, support vector machines, decision trees, and K-nearest neighbor algorithms [8]. In contrast, post-hoc methods are employed alongside AI techniques in a post-prediction setting to explain the (otherwise non interpretable) AI predictions and unveil nonlinear mappings within complex datasets. In this manuscript, we will focus on post-hoc ”local” methods that provide an explanation for each AI prediction separately, as opposed to post-hoc ”global” methods that derive explanations for the decision-making of the AI for the entire dataset [9]. One notable post-hoc local technique is the local interpretable model-agnostic explanations (LIME), which explains the network’s predictions by building simple interpretable models that approximate the deep network locally, i.e., in the close neighborhood of the prediction of interest [10]. Post-hoc techniques encompass model-specific approaches designed to address specific nonlinear model behaviors and model-agnostic approaches to investigate data complexity [3, 8]. In computer vision, model-agnostic techniques such as LIME and perturbation-based methods find extensive application, while model-specific methods include feature relevance, condition-based explanation, and rule-based learning [11, 12, 8]. Similarly, in medical imaging, XAI methods predominantly consists of attribution and perturbation techniques [13]. Attribution techniques identify the important features for a given prediction by assigning relevance scores to features of a given input. Perturbation techniques assess the sensitivity of an AI prediction to certain input features [14], by systematically perturbing sub-groups of the input data [13, 15]. Among the XAI methods used in medical imaging, GRAD-CAM emerges as one of the most frequently employed techniques [15].
Despite the benefits and potential of XAI methods in advancing model trust, model fine-tuning and learning new science ([16]), recent studies have shown that they suffer from important pitfalls ([14, 17]). For instance, when using the rule of the method Layer-wise Relevance Propagation (LRP, an attribution method; [18]), explanations have been shown to be smooth and comprehensible but of lower accuracy/faithfulness to the deep network decision making ([14]). Reportedly, this is because the rule considers only the positive preactivations in the network ([14, 17]). Recent studies have suggested LRP modifications that combine various LRP rules to produce both faithful and smooth results ([19]). As a second example, SHapley Additive exPlnations (SHAP;[20]), an attribution method based on game theory principles, computes relevance scores called Shapley values for input features, for a given prediction. The primary challenge of implementing SHAP lies in accurately calculating the Shapley values by sampling all theoretically possible combinations of features (i.e., in line with game theory principles), due to constraints raised by computing resources. Several studies have utilized adapted Monte Carlo methods within the permutation space to achieve more precise sampling of Shapley values [21]. Moreover, efforts have been made to improve results by minimizing the weighted least square error of Shapley values estimation, leveraging learned parametric functions, as described in studies of Kernel-Shap and Fast-Shap [22]. Similarly, SVARM ([23]) employs stratified sampling by segmenting a dataset into sub-groups or strata and sub-sampling from each input instance. By combining mean estimates from all strata, this technique achieves a more precise estimation of the Shapley values and effectively mitigates the variance. Despite these efforts to optimize already established XAI methods, no existing XAI method has been proven to be consistently superior when considering various performance scores and properties that an ideal method should exhibit (e.g., faithfulness to the network decision making process, comprehensibility, robustness, localization; [24]). As such, a more holistic approach (implementing many methods) is typically preferred ([16]). However, even when explaining the same deep network and the same prediction, different XAI methods yield highly different explanations (see Fig. 1), increasing uncertainty about the decision making process of deep networks and lowering trust. The sub-optimal results of individual XAI methods combined with the high inter-method variability constitute one of the biggest challenges in the field of XAI currently.
@@
In this study, for the first time, an ”explanation optimizer” is proposed to derive an optimal explanation for deep classifiers in computer vision tasks, based on specific pre-defined performance metrics (see Fig. 1), thus, solving the enigma of explainability. The optimizer reconstructs a singular explanation by integrating the outputs of existing XAI methods. This approach ensures that the resulting explanation both accurately captures the essence of the network’s decision-making process (honoring faithfulness) and maximizes comprehensibility (minimizing complexity). The framework ultimately provides a unique optimal explanation with the same or higher resolution as the network’s input (see Fig. 2). We note that providing high resolution explanations is useful for more accurate identification of significant features, particularly crucial in medical applications [9].
The figure depicts the traditional local explanation process (implementing multiple XAI methods) for a three-dimensional binary classification example regarding the existence or absence of PCS (Paracingulate Sulcus Detected). This is one of the two case studies we consider in this paper. Note that the high variability in the local explanations that arises from different XAI methods increases uncertainty and lowers trust in explainability among experts. Therefore, an explanation optimizer is proposed to derive the optimal explanation for the pre-trained Deep Classifier in the computer vision task of interest.
To assess the effectiveness of our framework, we conducted experiments on both multi-class and binary classification tasks in two- and three-dimensional spaces, respectively, focusing on automobiles-animals and neuroscience imaging domains. For the three-dimensional task, we utilized a comprehensive dataset consisting of 596 subjects with 3D brain volumes, sourced from the TOP-OSLO study [25]. This dataset was employed for binary classification, specifically exploring the presence or absence of the paracingulate sulcus (PCS). The PCS is a variable secondary sulcus, which is associated with cognitive performance and hallucinations in patients with schizophrenia [26],[27]. For the two-dimensional multi-class task, we employed the CIFAR-10 dataset [28] a widely-used benchmark dataset, containing approximately 60,000 images across 10 different classes of animals and vehicles. In both experiments, we partitioned the datasets into training (60%), validation (20%), and testing (20%) subsets to ensure robust evaluation of our framework’s performance.
@@
The explanation optimizers, employed in the 3D and 2D experiments, are presented in Fig. 2a. and b, respectively. In both cases, the optimizer is initiated with a layer of local explanations from various already-established XAI methods. A weighted average of these explanations is computed. The individual local explanations and the weighted average are then concatenated and fed into a deep learning network (non-linear structure) to reconstruct a unique, optimal explanation (see Methods section for more details). After the optimal explanation is derived, a high-resolution version is extracted by using an Up-Sampling layer (see Fig. 2a.b). Finally, for training, the cost function evaluates i) the similarity between the weighted average and low-resolution prediction, ii) the faithfulness score of the explanation as quantified in [24], and iii) the complexity score of the explanation as quantified in [24] (Fig. 2a.b). The similarity score above (i) is used in the loss function to constrain the optimization solution and to accelerate the training convergence. Additional details can be found in the Methods section.
The inspiration behind the proposed explanation optimizer is rooted in Herbert Simon’s seminal work on ”bounded rationality” and ”satisficing” decision-making [29, 30]. Bounded rationality acknowledges that individual rationality is constrained by the heuristics and biases humans employ to explain the world. Heuristics are cognitive rules individuals use in their daily lives to make routine decisions, while cognitive biases are specific beliefs or perspectives supported by evidence and reasoning. Thus, ”bounded rationality” recognizes the limitations of human knowledge, time, and computational power. Simon introduced the term “satisficing” in 1956, representing the concept where individuals seek solutions or accept choices that are ”good enough” for their purposes, even if they are not fully optimized [31]. This concept provides an alternative basis for decision-making modeling, complementing the traditional view of decision-making as optimization. Simon’s ideas of bounded rationality and satisficing decision-making have gained widespread acceptance, influencing research across the social sciences. In our application, we leverage bounded rationality by integrating established knowledge from different XAI methods that may or may not be biased/bounded in different ways. Additionally, our proposed cost function serves as a criterion for ”satisficing” explanations in AI tasks.

### v2

@@ Abstract
The accelerated progress of artificial intelligence (AI) has popularized deep learning models across various domains, yet their inherent opacity poses challenges, particularly in critical fields like healthcare, medicine, and the geosciences. Explainable AI (XAI) has emerged to shed light on these ’black box’ models, aiding in deciphering their decision-making processes. However, different XAI methods often produce significantly different explanations, leading to high inter-method variability that increases uncertainty and undermines trust in deep networks’ predictions. In this study, we address this challenge by introducing a novel framework designed to enhance the explainability of deep networks through a dual focus on maximizing both accuracy and comprehensibility in the explanations. Our framework integrates outputs from multiple established XAI methods and leverages a non-linear neural network model, termed the ’explanation optimizer,’ to construct a unified, optimal explanation. The optimizer uses two primary metrics—faithfulness and complexity—to evaluate the quality of the explanations. Faithfulness measures the accuracy with which the explanation reflects the network’s decision-making, while complexity assesses the comprehensibility of the explanation. By balancing these metrics, the optimizer provides explanations that are both accurate and accessible, addressing a central limitation in current XAI methods. Through experiments on multi-class and binary classification tasks in both 2D object and 3D neuroscience imaging, we validate the efficacy of our approach. Our explanation optimizer achieved superior faithfulness scores, averaging 155% and 63% higher than the best-performing individual XAI methods in the 3D and 2D applications, respectively, while also reducing complexity to enhance comprehensibility. These results demonstrate that optimal explanations based on specific quality criteria are achievable, offering a solution to the issue of inter-method variability in the current XAI literature and supporting more trustworthy deep network predictions.
@@
[…]
We validate our framework in multi-class and binary classification tasks in 2D object and 3D neuroscience imaging, aiming to address inter-method variability and boost trust in AI predictions.
Our approach achieves faithfulness scores 63% and 155% higher than the best XAI method in the 2D and 3D applications, respectively, while at the same time reducing complexity and increasing comprehensibility.
We propose a generalized version of the framework for any explanation task in a multi-dimensional space.
1 Introduction
@@
In the last decade, deep learning models have led to transformative changes in various domains, including geosciences (Ham et al., 2019; Reichstein et al., 2019), automotive engineering, healthcare (Esteva et al., 2019; Samek et al., 2021), and biology (Greener et al., 2022; Mamalakis et al., 2024a). However, the inherent opacity of these sophisticated models, often likened to "black boxes," presents formidable challenges in understanding their decision-making processes. In response to this challenge, the field of eXplainable Artificial Intelligence (XAI) has emerged as a pivotal area of research dedicated to explaining these black-box models and elucidating the rationale behind their predictions. Over the years, a multitude of XAI methods have been developed, all aimed at shedding light on how artificial intelligence (AI) arrives at specific predictions.
Among others, XAI consists of methods that try to explain complex, nonlinear models by approximating them locally with simple interpretable models, such as local interpretable model-agnostic explanations (LIME ; Ribeiro et al. (2016)) and SHapley Additive exPlanations (SHAP ; Lundberg and Lee (2017)). In medical imaging, specifically, XAI methods often include attribution techniques like SHAP and Layer-wise Relevance Propagation (LRP ; Bach et al. (2015)) or sensitivity techniques, with GRAD-CAM being a prominent example (Singh et al., 2020). Despite their potential to enhance model trust and learning, XAI methods face significant challenges. For instance, LRP produces smooth results but less accurate explanations due to its focus on positive preactivations, and SHAP struggles with computational intensity in calculating Shapley values (Lundberg and Lee (2017)). Although adaptations such as Monte Carlo methods and stratified sampling (e.g., SVARM) have improved estimation precision (Kolpaczki et al., 2024), it has been established that no single XAI method consistently excels across all prediction tasks (Mamalakis et al., 2022), leading to high variability and uncertainty in deep learning explanations (Hedström et al., 2023).
@@
In this study, for the first time, an "explanation optimizer" is proposed to derive an optimal explanation for deep classifiers in computer vision tasks, based on specific pre-defined performance metrics (see Fig. 1), thus, solving the enigma of explainability. The optimizer, a non-linear neural network model, reconstructs a unified explanation by integrating outputs from various XAI methods. This model is guided by two critical metrics: faithfulness, which measures how accurately the explanation reflects the network’s decision-making process, and complexity, which assesses how comprehensible the explanation is. By balancing these scores, the optimizer aims to produce explanations that are both true to the model’s internal logic and accessible in their clarity. The framework ultimately provides a unique optimal explanation with the same or higher resolution as the network’s input (see Fig. 2). We note that providing high resolution explanations is useful for more accurate identification of significant features, particularly crucial in medical applications (Longo et al., 2024).
To assess the effectiveness of our framework, we conducted experiments on both multi-class and binary classification tasks in two- and three-dimensional spaces, respectively, focusing on automobiles-animals and neuroscience imaging domains.
The proposed explanation optimizer is inspired by Herbert Simon’s concepts of "bounded rationality" and "satisficing" decision-making (Simon, 1955, 1990). "Bounded rationality" reflects the limitations of human knowledge, time, and computational capacity, framing rationality as constrained by heuristics and cognitive biases that aid in efficient yet simplified decision-making. Simon’s notion of "satisficing," introduced in 1956, emphasizes finding solutions that are "good enough" rather than fully optimized (Simon, 1956). In our framework, we apply these principles by integrating complementary insights from various XAI methods to achieve explanations that balance high faithfulness with manageable complexity. Our novel cost function supports the idea of "satisficing" explanations, targeting optimal interpretability (as quantified by specific metrics) for specific computer vision tasks across multiple domains, including automobiles, animals, and neuroscience. Furthermore, we leverage the concept of ’bounded rationality’ by integrating established knowledge from different XAI methods as inputs to the optimizer, recognizing that these methods may or may not be biased or bounded in different ways.
To encapsulate, in this study, we propose a framework aimed at providing optimal explanations for pre-trained deep networks in specific computer vision tasks. Characterized by high faithfulness and low complexity, this framework ensures accuracy and comprehension in the provided explanations. Additionally, we develop a separate framework designed to deliver explanations in high-resolution spaces, enhancing our ability to capture and understand significant feature contributions effectively. Our framework is tested in applications across diverse domains, including automobiles, animals, and neuroscience, encompassing both 2D and 3D multi-classification and binary classification tasks. Finally, we formulate a generalized framework for optimizing explanations for M-dimensional inputs, extending the utility of our approach beyond specific computer vision tasks to offer a versatile solution for interpreting complex data across various dimensions and domains. Our results suggest that optimal explanations based on specific criteria are indeed derivable and address the issue of inter-method variability in the current XAI literature.
2 Related Work
Explainable artificial intelligence can be broadly classified into two methodological approaches: inherently interpretable and post-hoc. Interpretable methods prioritize models endowed with properties such as simulatability, decomposability, and transparency, making them inherently comprehensible. Typically, these include linear techniques like Bayesian classifiers, support vector machines, decision trees, and K-nearest neighbor algorithms (van der Velden, 2023). In contrast, post-hoc methods are employed alongside AI techniques in a post-prediction setting to explain the (otherwise non interpretable) AI predictions and unveil nonlinear mappings within complex datasets. In this manuscript, we will focus on post-hoc "local" methods that provide an explanation for each AI prediction separately, as opposed to post-hoc "global" methods that derive explanations for the decision-making of the AI for the entire dataset (Longo et al., 2024). One notable post-hoc local technique is the local interpretable model-agnostic explanations (LIME), which explains the network’s predictions by building simple interpretable models that approximate the deep network locally, i.e., in the close neighborhood of the prediction of interest (Ribeiro et al., 2016). Post-hoc techniques encompass model-specific approaches designed to address specific nonlinear model behaviors and model-agnostic approaches to investigate data complexity (Samek et al., 2021; van der Velden, 2023). In computer vision, model-agnostic techniques such as LIME and perturbation-based methods find extensive application, while model-specific methods include feature relevance, condition-based explanation, and rule-based learning (Ribeiro et al., 2016; Rajani et al., 2019; van der Velden, 2023). Similarly, in medical imaging, XAI methods predominantly consists of attribution and perturbation techniques (Tjoa and Guan, 2020). Attribution techniques identify the important features for a given prediction by assigning relevance scores to features of a given input. Perturbation techniques assess the sensitivity of an AI prediction to certain input features (Mamalakis et al., 2022), by systematically perturbing sub-groups of the input data (Tjoa and Guan, 2020; Singh et al., 2020). Among the XAI methods used in medical imaging, GRAD-CAM emerges as one of the most frequently employed techniques (Singh et al., 2020).
Despite the benefits and potential of XAI methods in advancing model trust, model fine-tuning and learning new science (McGovern et al., 2019), recent studies have shown that they suffer from important pitfalls (Mamalakis et al., 2022; Dikshit et al., 2024). For instance, when using the rule of the method Layer-wise Relevance Propagation (LRP, an attribution method; Bach et al. (2015)), explanations have been shown to be smooth and comprehensible but of lower accuracy/faithfulness to the deep network decision making (Mamalakis et al., 2022). Reportedly, this is because the rule considers only the positive preactivations in the network (Mamalakis et al., 2022; Dikshit et al., 2024). Recent studies have suggested LRP modifications that combine various LRP rules to produce both faithful and smooth results (Kohlbrenner et al., 2020). As a second example, SHapley Additive exPlanations (SHAP;Lundberg and Lee (2017)), an attribution method based on game theory principles, computes relevance scores called Shapley values for input features, for a given prediction. The primary challenge of implementing SHAP lies in accurately calculating the Shapley values by sampling all theoretically possible combinations of features (i.e., in line with game theory principles), due to constraints raised by computing resources. Several studies have utilized adapted Monte Carlo methods within the permutation space to achieve more precise sampling of Shapley values (Mitchell et al., 2022). Moreover, efforts have been made to improve results by minimizing the weighted least square error of Shapley values estimation, leveraging learned parametric functions, as described in studies of Kernel-Shap and Fast-Shap (Jethani et al., 2022). Similarly, SVARM (Kolpaczki et al., 2024) employs stratified sampling by segmenting a dataset into sub-groups or strata and sub-sampling from each input instance. By combining mean estimates from all strata, this technique achieves a more precise estimation of the Shapley values and effectively mitigates the variance. Despite these efforts to optimize already established XAI methods, no existing XAI method has been proven to be consistently superior when considering various performance scores and properties that an ideal method should exhibit (e.g., faithfulness to the network decision making process, comprehensibility, robustness, localization; Hedström et al. (2023)). As such, a more holistic approach (implementing many methods) is typically preferred (Mamalakis et al., 2022). However, even when explaining the same deep network and the same prediction, different XAI methods yield highly different explanations (see Fig. 1), increasing uncertainty about the decision making process of deep networks and lowering trust. The sub-optimal results of individual XAI methods coupled with the high inter-method variability constitute one of the biggest challenges in the field of XAI currently and is the focus of this work.
3 Methodology
3.1 Evaluation metrics for the explanation
A crucial aspect of this study lies in evaluating "how accurate and comprehensive is an explanation?" To derive a useful explanation, two primary scores play a pivotal role: faithfulness and complexity. An intuitive way to assess the quality of an explanation is by measuring its ability to accurately capture how the predictive model reacts to random perturbations (Yeh et al., 2019). For a deep neural network , and input features , the feature importance scores (also known as "attribution scores") are derived in a way such that when we set particular input features to a baseline value , the change in the network’s output should be proportional to the sum of attribution scores of the perturbed features . We quantify this by using the Pearson correlation between the sum of the attributions of and the difference in the output when setting those features to a reference baseline (Bhatt et al., 2020). Thus, we define the faithfulness of an explanation method as:
where is a subset of indices (), is a sub-vector of an input ( and the unchanged features of image). The total number of features, which partition an image is . We denote as the changed features, and the unchanged features of image.
If for an image , explanation highlights all features, then it may be less comprehensible and more complex than needed (especially for large ). It is important to compute the level of complexity, as an efficient explanation has to be maximally comprehensible (Bhatt et al., 2020). If is a valid probability distribution and the is the fractional contribution of feature to the total magnitude of the attribution, then we define the complexity of the explanation for the network as:
where:
In order to evaluate these two explainability metrics, we used the software developed by (Hedström et al., 2023). This software package is a comprehensive toolkit that collects, organizes, and evaluates a wide range of performance metrics, proposed for XAI methods. We note that we used a zero baseline (’black’; = ), and 70 random perturbations to calculate the Faithfulness score.
3.2 Overview of the proposed framework
@@
The proposed framework entails an optimization approach aimed at attaining the highest faithfulness and lowest complexity of an explanation tailored to a specific deep learning network and computer vision task. The framework was applied in 3D and 2D experiments, and the architectures of the optimizers are presented in Fig. 2 and in Supplementary material Fig. 1, respectively. In both cases, the optimizer is initiated with a layer of local explanations from various already-established XAI methods. A weighted average of these explanations is computed (’Weighted Average’). The individual local explanations and the weighted average are then concatenated and fed into a deep learning network (’Non-linear structure’) to reconstruct a unique, optimal explanation. After the optimal explanation is derived, a high-resolution version is extracted by using an up-sampling (’Up-Sampling’) layer (see Fig. 2). Finally, for training, the cost function evaluates i) the similarity between the weighted average and low-resolution prediction, ii) the faithfulness score of the explanation as quantified in (Hedström et al., 2023), and iii) the complexity score of the explanation as quantified in (Hedström et al., 2023) (Fig. 2). The similarity score above (i) is used in the loss function to constrain the optimization solution and to accelerate the training convergence.
Key components of the framework include the utilization of the baseline XAI methods, the ’Weighted Average’ explanation, the ’Non-linear structure’, and the ’Up-Sampling’ unit (Fig. 2). The assigned weights of the ’Weighted Average’ explanation are computed from the equation below:
where and represent fixed values ranging from 0 to 1, signifying the contributions of faithfulness and complexity, respectively, to the final weighted average. In this study, we analyzed predictions and explanations for ten randomly selected prediction instances from the validation cohort. This selection is made solely for computational efficiency. During training, for each batch, these 10 selected instances serve as a proxy to approximate the relative weights, allowing for the calculation of the weighted average of the explanations. We assessed the faithfulness and complexity scores for each of the explanation methods and for each prediction instance. Subsequently, we calculated the average faithfulness and complexity scores across the ten prediction instances and determined the final assigned weights using eq. 4 with . The average values of , and the assigned weights were normalized to range from 0 to 1.
The explanations generated by the baseline XAI methods are fed into a non-linear network consisting of an encoder, bottleneck, and decoder architecture (see later sections for more details), with the objective of reconstructing a unique explanation. Following this, the reconstructed explanation undergoes sequential processing through an up-sampling deconvolution layer to double resolution. In our case, the resolutions of the 3D inputs in the application were , while for the 2D images, they were . The final output of our optimizing framework consists of an up-sampled higher-resolution explanation based on the input ( in the 3D application and in the 2D application) and an explanation with the same resolution as the input (, and , respectively). The up-sampling approach is motivated by the aim to enhance the detail of the explanation. We note that particularly in neuroscience and medical applications, a higher resolution is closely linked with increased trust of the community to AI procedures, making the provision of explanations in both higher and intial resolutions highly desirable (Moerel et al., 2021; Turner and De Haan, 2017; Feinberg et al., 2023; Markello et al., 2022).
where represents the derived explanation by our optimizer, denotes the ’Weighted Average’ explanation, indicates the average of , signifies the variance of , represents the covariance of and , and and are two parameters utilized to stabilize the division with a weak denominator (Wang et al. (2004)). refers to the similarity score between the ’Weighted Average’ explanation and the down-sampled predicted explanation (low resolution), while signifies the similarity score between a tri-linear or bicubic (in 3D and 2D space respectively) up-sampled ’Weighted Average’ explanation and the derived high-resolution explanation. The values of and were fixed at 0.5 each. The total loss function was given by:
Following an ablation study exploring various combinations of the values for and in the above loss function, we determined that setting and is optimal. This selection ensures a balanced optimization approach that effectively addresses the unique challenge of deriving an optimal explanation.
3.3 The baseline XAI methods
For this study, we employed eight established XAI methods (): Saliency (Lundberg and Lee, 2017), DeepLift (Lundberg and Lee, 2017), Kernel Shap (Lundberg and Lee, 2017), DeepLift Shap (Lundberg and Lee, 2017), Integrated Gradients (Sundararajan et al., 2017), Guided Backpropagation (Springenberg et al., 2015), Guided GradCam (Selvaraju et al., 2019) and Gradient Shap. GradientShap makes an assumption that the input features are independent and that the explanation model is linear, meaning that the explanations are modeled through the additive composition of feature effects. Under these assumptions, SHAP values (Lundberg and Lee, 2017) can be approximated as the expectation of gradients computed for randomly generated samples from the input dataset. Gaussian noise is added to each input, creating different baselines for computing the SHAP values.
Extending the implementation of these explainability approaches from two-dimensional (2D) to three-dimensional (3D) space was necessary. We utilized the Python software developed by (Hedström et al., 2023) as the underlying libraries for our implementation. Further implementation details can be found in the public GitHub repository associated with this study. It’s important to note that all explanations were modified to focus exclusively on positive attributions. Any negative attribution in the original explanations were replaced with a zero value. As part of future work, we plan to incorporate negative attributions into the reconstruction process of the explanation optimizer to enhance interpretability.
3.4 Deep learning architectures
We utilized a Resnet-18 and a ResNet-50 network architectures (He et al., 2015) for the classification tasks in the three-dimensional (3D) and two-dimensional (2D) applications, respectively.
With regards to the explanation optimizer, for the 3D application, we employed the Swin-UNETR architecture (Hatamizadeh et al., 2022) as our non-linear reconstruction network, featuring a 24-feature size (Fig. 2). For the 2D explanation optimization, we utilized a U-NET encoder-decoder architecture (Ronneberger et al., 2015); see Fig. 1 in the Supplementary material).
For the sake of further investigation (see subsection 4.5.’Why utilizing a non-linear optimizer structure over linear alternatives?’), we explored a linear implementation in the three-dimensional application. Here, we employed visual transformer encoders (Dosovitskiy et al., 2020) to estimate weights for each of the explainable baseline methods (refer to Supplementary material subsection: ’Deep learning architectures: The linear implementation of the proposed network’). Subsequently, these estimated weights were utilized in a weighted average to extract the linearly optimized explanation.
4 Experiments and Results
4.1 The experiments in 3D and 2D applications
In the 3D application, our focus is on a binary classification task concerning the presence or absence of the paracingulate sulcus (PCS) in the brain. Leveraging a comprehensive cohort of 629 subjects from the TOP-OSLO study (Mørch-Johnsen et al., 2018). We utilized data from 596 patients after excluding 33 cases due to pre-processing errors in 3D topology extraction and inadequate PCS detection levels. For the 3D classification task, we utilize white-grey surface matter images extracted from T1-weighted MRI scans of patient brains. Our image preprocessing follows the outlined steps in (Mamalakis et al., 2024b). For the 2D application, we utilize the publicly available CIFAR-10 dataset (https://paperswithcode.com/dataset/cifar-10). This dataset, a subset of the Tiny Images dataset, consists of 60,000 colored images with dimensions of . Each image depicts and corresponds to one of 10 mutually exclusive labels (i.e., classes), such as airplane, automobile, bird, cat, deer, dog, frog, horse, ship, and truck.
4.2 Data statement for the 3D TOP-OSLO dataset
The data supporting the findings of the present study are hosted in the repository at NORMENT/Oslo University Hospital. Restrictions apply to the availability of data and they are thereby not publicly available. Data can be made available under reasonable request and with permission of NORMENT/Oslo University Hospital, in accordance with the ethics agreements/research participants consent.
4.3 Implementation details
For both the 2D and 3D classification tasks, after randomly shuffling the data, each dataset underwent splitting into training, validation, and testing sets, with proportions of , , and of the total number of images, respectively. Sparse categorical cross-entropy was employed as the cost function, while the Adam algorithm (Kingma and Ba, 2014) was utilized to optimize the loss function. Initially, the learning rate remained fixed for the first 300 epochs, followed by a decrease of every 100 epochs. To prevent overfitting, an early stopping criterion of 100 consecutive epochs was enforced, with a maximum of 500 epochs used for the input modalities in the corresponding 2D and 3D applications. Data augmentation techniques were applied, including rotation (around the center of the image by a random angle from the range of ), width shift (up to pixels), height shift (up to pixels), and Zero phase Component Analysis (ZCA;Bell and Sejnowski (1997)) whitening (adding noise to each image). Hyperparameter tuning was conducted for both applications using four different values of initial learning rates: , , , and .
The same fixed-step learning rate and optimization algorithm, namely the Adam algorithm (Kingma and Ba, 2014), were employed to optimize the loss function in the explainability task. However, no data augmentation techniques were utilized to train the optimizer. The cost function used was that of eq. 7. For the 2D task, the maximum number of epochs was set at 150, with an early stopping criterion of 10 epochs applied after passing the first 90 epochs. For the 3D task, the maximum number of epochs was set at 500, with an early stopping criterion of 120 epochs applied after the initial 300 epochs. Once again hyperparameter tuning was conducted for both applications using four different values of initial learning rates: , , , and .
The code developed in this study is written in the Python programming language using pyTorch (Python) libraries. For the training and testing of deep learning networks, we have used an NVIDIA cluster with GPUs and GB RAM memory. We used an NVIDIA A100-SXM-80GB GPU, and the code will be publicly accessible through github.
4.4 Results of classification training and optimization of explanations
The deep neural networks were trained for binary and multi-label classification tasks, in three and two dimensional spaces. The sensitivity of the results for the different learning rates is illustrated with light blue shading in Fig. 3 and in Supplementary material Fig. 3. Results show the minimum, median and maximum loss and accuracy values that the deep networks achieved in the training and validation datasets for the 3D and 2D applications, respectively. In the 3D binary classification task, the network achieved approximately 75.0% validation accuracy and around 70.0% training accuracy (see Fig. 3). For the 2D multi-classification task, the highest validation accuracy exceeded 80.0%, with training accuracy surpassing 75.0% (see Supplementary material Fig. 3). Fig. 3 and Supplementary material Fig. 3 demonstrate the effective training of Resnet-18 and ResNet-50 models on Top-OSLO and CIFAR-10 datasets, respectively. The training and validation performance is shown to be fairly similar, indicating a low risk of overfitting. We note that the optimal learning rate for the three-dimensional application was found to be , while for the two-dimensional application, it was .
Next, Fig. 4 and Supplementary material Fig. 4 present the loss function during training of the explanation optimizer for each class in the 3D and 2D tasks. The light blue shading in Fig. 4a. and Supplementary material Fig. 4 indicates the min and max loss values, while the solid currve represents the median loss across different learning rates. The best performance for the optimizer was achieved in the three-dimensional application for a learning rate of for the presence of PCS, for the absence of PCS, and for the two-dimensional scenario.